OpenStudy (anonymous):

5x-10x=-40 solve for y

OpenStudy (anonymous):

Um, should one of those variables be "y"??

OpenStudy (anonymous):

sorry it is this 5x-10y=-40

OpenStudy (anonymous):

Lol. -10y=-5x-40 y= 1/2 x + 4 And thats your answer. Get it?

OpenStudy (anonymous):

thanks so much how did you get that answer

OpenStudy (anonymous):

When solving for y, you want everything on one side. 5x- 5x -10y= -5x-40 -->>subtract 5x from both sides -10y/-10= -5x/-10 - 40/-10 -->> divide eveything by -10 Is it clearer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

great! you're welcome

OpenStudy (anonymous):

here is another one 2x+3/4 =x/4 solve

OpenStudy (anonymous):

which is "y"?? lol

OpenStudy (anonymous):

none it is not solving for y

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Oh, duuuh. Lol. 4( 2x+3/4)= x -->> multiply each side by 4 8x+3= x 3= -7x -->> combine like terms by subtracting x and 8x. So 1-8= -7 -3/7= x --->> divide each side by -7. Here's your answer........hopefully. Im not sure about this one.

OpenStudy (anonymous):

Thats sooo wrong. it was supposed to be: 8x+12=x 12=-7x -12/7 = x

OpenStudy (anonymous):

you have cross mutiply

OpenStudy (anonymous):

it is not sppouse to be a freaction

OpenStudy (anonymous):

o__0 Cross multiply?? I dont see how. Im sorry I cant help. Mayb you can turn the fraction into a decimal??

OpenStudy (anonymous):

no it is 1 single number here are the answer a.x=-3 b.x=-1 c.x=1 d.x=3

OpenStudy (anonymous):

What?!? Who is your teacher, this is ridiculous! This is why i hate math. 1) -3/7 was correct. Ignore the second part. 2) I am so so sooo sorry I can help.

OpenStudy (anonymous):

cant**